Find the gradient vector field of f.

f(x,y,z)=x^2ye^{^{y/z}}

I need Matlab Code Also
Answer #1

Matlab Code:

syms x y z
f = x^2*y*exp(y/z)
gradient(f, [x, y, z])

syms x y z f = x^2*y*exp (y/z) gradient (f, (x, y, z])

Solution:

>> graddd

f =

x^2*y*exp(y/z)


ans =

2*x*y*exp(y/z)
x^2*exp(y/z) + (x^2*y*exp(y/z))/z
-(x^2*y^2*exp(y/z))/z^2

>>

Command Window >> graddd f = x^2*y*exp(y/z) ans = 2*x*y*exp(y/z) x^2*exp (v/z) + (x^2*y*exp (v/z))/z -(x^2y^2*exp(y/z))/2^2 f

Hand calculations:

zz -੦) 2 - z e xe fet gradient vector > = ? (%) ਹਰਮੰ = ਦੋ ਸz > ਜਦ ( ++ 6 = = = = ਵਿਚ। Z= (ਦੋ ਦੋ :=(ZVWਸf

-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-

KINDLY UPVOTE THE ANSWER IF YOU LIKED IT, IT WILL MAKE MY DAY ?

FOR ANY QUERIES COMMENT BELOW.

-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-_-